OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 2x - 3y - z = -1 x - 4y + 3z = -11 2x + 9y + 3z = -13 A. -11y + 5z = 21 y + 9z = 9 B. 5y - 7z = 21 17y - 3z = 9 C. 7x - 13y = -14 8x + 0y = -16 D. -5x +5y = -8 -4x + 18z = -10

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@texaschic101 can you help me with this?

OpenStudy (texaschic101):

x - 4y + 3z = -11 x = 4y - 3z - 11 2x - 3y - z = -1 2(4y - 3z - 11) - 3y - z = -1 8y - 6z - 22 - 3y - z = -1 5y - 7z = 21 <==== here is one new equation 2x + 9y + 3z = -13 2(4y - 3z - 11) + 9y + 3z = -13 8y - 6z - 22 + 9y + 3z = -13 17y - 3z = 9 <==== here is the other new equation answer is B

OpenStudy (texaschic101):

I just isolated x in the second equation and subbed that in for x in the other equations, and I just simplified the equations